OpenStudy (anonymous):

A cylindrical rain barrel has a radius of 2 feet and holds a total of 30 cubic feet of water. How tall is the rain barrel? Use 3.14 for pi. Round your answer to the nearest hundredth. 1.58 ft 2.39 ft 3.57 ft 4.78 ft

OpenStudy (anonymous):

@saseal @ali2x2

OpenStudy (anonymous):

@heretohelpalways

OpenStudy (anonymous):

work with the cylinder equation

OpenStudy (anonymous):

30 = (3.14)(2^2)h 30 = (3.14)(4)h 30 = 12.56h 30 / 12.56 = ?